# A Circuit-QED Lattice System with Flexible Connectivity and Gapped Flat Bands for Photon-Mediated Spin Models

### v1

@@ Abstract
Quantum spin models are ubiquitous in solid-state physics, but classical simulation of them remains extremely challenging. Experimental testbed systems with a variety of spin-spin interactions and measurement channels are therefore needed. One promising potential route to such testbeds is provided by microwave-photon-mediated interactions between superconducting qubits, where native strong light-matter coupling enables significant interactions even for virtual-photon-mediated processes. In this approach, the spin-model connectivity is set by the photonic mode structure, rather than the spatial structure of the qubit. Lattices of coplanar-waveguide (CPW) resonators have been demonstrated to allow extremely flexible connectivities and can therefore host a huge variety of photon-mediated spin models. However, large-scale CPW lattices have never before been successfully combined with superconducting qubits. Here we present the first such device featuring a quasi-1D CPW lattice with a non-trivial band structure and multiple transmon qubits. We demonstrate that superconducting-qubit readout and diagnostic techniques can be generalized to this highly multimode environment and observe the effective qubit-qubit interaction mediated by the bands of the resonator lattice. This device completes the toolkit needed to realize CPW lattices with qubits in one or two Euclidean dimensions, or negatively-curved hyperbolic space, and paves the way to driven-dissipative spin models with a large variety of connectivities.
@@ I Introduction
Quantum spin models underlie a wide variety of solid-state phenomena from magnetism, to spin glasses, to parts of the phase diagram of high-temperature superconductors [1, 2, 3]. Efficient classical simulation algorithms exist in special types of connectivity, but the general phenomenology of these systems remains an outstanding grand challenge, particularly when a spin Hamiltonian is accompanied by drive and dissipation. Implementing these models in synthetic quantum systems, which offer access to novel local or dynamical measurements, holds the potential to provide new insights into this challenge. Here we present an experimental platform, using superconducting qubits coupled to microwave resonator arrays, capable of implementing spin-spin interactions with a large variety of connectivities, spanning from simple 1D chains, to 2D, to positively-curved spherical spin models, and even to hyperbolic spin models [4, 5].
@@
The crucial ingredient underlying this flexibility is the use of microwave photons to mediate interactions between effective spin-$1/2$ degrees of freedom formed by qubits. Direct qubit-qubit coupling inherently gives rise to a spin-spin interaction which follows the geometry of the device. By contrast, using a photon to mediate coupling allows the geometry of the spin-spin interaction to be decoupled from the physical locations of the qubits, and in these cases, the limits on achievable interactions are set by the ability to control and modify the propagation of the mediating photon. In the simplest case, a single-mode cavity mediates an all-to-all interaction between qubits, leading to trivial connectivity. An interaction that generates a spin model with non-trivial connectivity requires the presence of multiple modes in a bandwidth set by the qubit-photon interaction strength [6, 5, 7, 8]. In such cases, the dispersion relation and spatial profiles of the photon modes determine the interaction profile.
@@
The circuit QED platform [9], featuring superconducting qubits coupled to microwave resonators, is an ideal host for such models due to the combination of two types of hardware properties. (i) Superconducting microwave resonators have very high quality factors and can be implemented in a wide variety of different ways, with highly variable form factors on chip. As a result of this flexibility, superconducting microwave resonator arrays have been shown to host linear, quadratic, and flat bands in both one and two dimensions [10, 11, 4, 12, 13, 14, 15, 16, 17], and non-Euclidean structures [18]. Flat bands are particularly important since they quench kinetic energy, permitting interaction-driven strongly correlated phases to emerge in many-body systems. Furthermore, beyond their intrinsic interest, many-body systems on manifolds with spatial curvature enable experimental access to exotic topological observables and can even challenge the basic framework for analyzing quantum systems. (ii) Superconducting qubits are frequency tunable, and natively exhibit strong light-matter coupling, which can easily be adjusted by orders of magnitude. The combination of these two properties gives unprecedented ability to tailor the spatial structure of strong qubit-qubit interactions, without the loss in the photon modes significantly damaging native qubit coherence times.
@@
Multimode cavity-QED models with the potential to host boson-mediated spin models with non-trivial connectivity have also been studied in atomic systems with multiple discrete modes [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]; or a continuum of modes from a waveguide [35, 36], a photonic crystal [6], or a BEC [37, 38, 39]; or with Raman coupling in a single-mode cavity [40]. However, none of these systems can achieve the large coupling strengths and the variety of connectivities possible in superconducting microwave circuits.
@@
Previous multimode QED and waveguide QED experiments with superconducting circuits have used modulated waveguides [41, 7], simple one-dimensional chains of resonators [42, 17, 14, 13, 15], or 2D Hofstadter lattices [43, 44]. In these cases, the attainable dispersion relations and spatial profiles of the photon modes have been limited by either the simplicity of the waveguide chain or by the rigid form factor of the 3D cavities. In contrast, coplanar-waveguide (CPW) resonators [45] can form resonator arrays [4, 11, 46, 10] with low-disorder [11] and highly flexible connectivities [18, 12]. It has been demonstrated that this technique can produce lattices in which the photons propagate in an effective negatively-curved hyperbolic space [18], and it has been shown theoretically that the photon-mediated qubit-qubit interaction that a hyperbolic CPW lattice would produce also follows the hyperbolic metric [5]. These lattice systems also have a strong tendency to produce flat bands that consist entirely of wave functions which flip sign on neighboring sites [18, 12], and which can therefore give rise to short-range frustrated magnetic interaction between qubits. However, all previous studies of coplanar-waveguide arrays have either been purely theoretical [4, 46, 12] or featured a resonator array without qubits [10, 11, 18].
@@
Here we present the first CPW lattice device featuring both a large-scale resonator array with non-trivial band structure and multiple superconducting qubits. Our device features a Euclidean quasi-one-dimensional microwave resonator lattice with a non-trivial unit cell, which gives rise to conventional quadratic bands, linearly dispersing bands, and flat bands with localized eigenstates. We incorporate three flux-tunable transmon qubits directly into the CPW lattice and show that the low-disorder microwave resonator array survives the addition of the qubits and their associated flux-bias control.
@@
We demonstrate that conventional superconducting-qubit readout and diagnostic techniques [9], though originally developed for single-mode systems, can be applied in this highly multimode environment, and can be used to observe the effective qubit-qubit interaction mediated by the bands of the resonator lattice. Additionally, we develop a novel in situ diagnostic technique, which leverages the presence of qubits, to measure the mode spectrum of the CPW lattice. Previous related techniques were either extremely time-consuming [11] or highly sensitive to the location of probe ports and unable to distinguish lattice modes from some types of parasitic modes induced by device packaging [18]. In contrast, our method is fast, shows very strong suppression of the effects of background modes, and does not require two-port transmission of all modes, enabling observation of localized modes, as well as delocalized ones. The device and measurement techniques presented here pave the way to CPW-lattice-based spin models with frustrated, 2D, and even hyperbolic connectivity.
@@ II Background
Traditional cavity QED describes an interaction between a single qubit/atom/spin and a single photon mode. Extensions to a continuum of photon modes have been studied in the context of emitters coupled to a waveguide, known as waveguide QED [35, 36]. CPW lattices produce a large number of discrete modes which approximate the continuous density of states of the bands of an infinite lattice. These systems therefore approximate the full continuum of waveguide QED. To simplify the language and highlight the commonalities between different ways of creating multimode systems, we refer to all systems in which an effective continuum of highly-coupled modes with a dispersion relation is relevant to forming photon-mediated interactions as (generalized) photonic crystal and waveguide QED systems. In the remainder of this section, we present a general framework for describing multimode QED devices of this kind with multiple qubits. Characteristics of these models which are specific to superconducting circuits and CPW resonators, in particular, will be discussed in Section III.
@@ II.1 Cavity and Waveguide QED
The simplest Hamiltonian for a cavity-qubit coupled system is known as the Jaynes-Cummings Hamiltonian [47, 48, 49]:
@@
In this work we consider systems where the waveguide is formed by a resonator network. We therefore present the general discrete-mode case of waveguide QED here, but extension to a full continuum is straightforward [50]. In this framework the qubit and cavity portions of the Hamiltonian become
@@
[…]
@@ II.2 Photon Bound States and Photon-Mediated Interactions
Photonic band gaps [51] are typically created through some form of periodic modulation of the medium. One type of device architecture involves periodic modulation of the index of refraction, for the case of optical photons [52], or the impedance [41, 7], for microwave photons, and generally has a full continuum of delocalized modes. Another class of devices instead uses periodic arrays of localized sites such as optical or microwave resonators [11, 4, 18, 17, 43, 44, 13, 14, 16, 53]. In these devices, the photonic modes are delocalized only by the coupling between the sites, and can be thought of as akin to tight-binding models of solids [1]. The device presented in this work is chosen to be of the latter kind because the tight localization of the mode in the individual resonators leads to separation between the hardware properties that set the on-site energy and the properties that control the coupling, or hopping, between sites.
@@
In the context of photonic band gaps, photon-mediated qubit-qubit interactions are often described as arising from the formation of photonic bound states [51, 6, 54]. These are localized modes with photonic character which emerge even if all of the photon modes are delocalized in the absence of coupling to the qubit. If the qubit is in the band gap, there is an intuitive picture of the formation of the bound state in which the qubit hybridizes with the photon modes and acquires a cloud of virtual photons surrounding it. In the conventional case, where the dispersion of the photon modes above the band edge is quadratic ($\omega_{k}=\omega_{edge}+\beta k^{2}$) and the coupling constants $g_{j,k}$ are approximately equal, this photon cloud is known to take the form of an exponentially localized envelope, $e^{-|x|/L}$ [51, 6, 54], where $|x|$ is the distance from the qubit, and where the localization length $L$ is determined by both the band curvature $\beta$, and the detuning between the qubit and the band edge. The simple picture of the photonic bound state being a cloud of virtual photons surrounding a qubit is valid only for a qubit well inside the band gap. However, an exact solution of similar form exists at any detuning from an ideal quadratic band edge and has been derived in Refs. [51, 41, 7, 54]. If two qubits are coupled to the band edge, then interactions between them can be thought of as arising from the tail of the photonic bound state localized around one qubit driving transitions in the other qubit. Hence, the spatial structure of the photonic bound state dictates the spatial structure of the interaction.
@@
Alternatively, the photon-mediated qubit-qubit interactions can be modeled directly as a photon-exchange process [6, 7, 5], as described in Section II.1. For arbitrary detunings, the frequency of the photonic bound state and the resulting interaction between the $j^{th}$ qubit and the $l^{th}$ qubit must be found self-consistently [7, 5, 54]. However, if both qubits are in the band gap, then second-order perturbation theory can be used to obtain a simple form valid for any dispersion
@@
where $\sigma_{j,-}$ and $\sigma_{j,+}$ indicate raising and lowering operators on the $j^{th}$ qubit, and $\Delta_{k}$ is the detuning between the qubits and mode $k$ (assumed to be approximately equal for the two qubits). A quadratic band edge in a one-dimensional set of modes gives a simple exponential interaction, but this can be changed by going to higher dimensions, non-quadratic dispersions [55], or non-plane-wave modes.
@@ III Photonic Crystals in Circuit QED
Since the form of the qubit-qubit interaction mediated by a photonic bound state is determined by the dispersion and mode functions of the band, creating spin models with a large variety of connectivities requires the ability to engineer diverse band structures with strong qubit-photon coupling. In this section we describe the hardware aspects of CPW lattices that make CPW lattices ideally suited to this task, as well as the measurement techniques native to this microwave-frequency implementation of waveguide QED. We emphasize that the methods developed for single-mode circuit QED [9] can be generalized to a full toolkit for design, characterization, and measurement of multimode waveguide-QED devices made with CPW resonators.

### v2

@@ Abstract
Quantum spin models are ubiquitous in solid-state physics, but classical simulation of them remains extremely challenging. Experimental testbed systems with a variety of spin-spin interactions and measurement channels are therefore needed. One promising potential route to such testbeds is provided by microwave-photon-mediated interactions between superconducting qubits, where native strong light-matter coupling enables significant interactions even for virtual-photon-mediated processes. In this approach, the spin-model connectivity is set by the photonic mode structure, rather than the spatial structure of the qubit. Lattices of coplanar-waveguide (CPW) resonators have been demonstrated to allow extremely flexible connectivities and can therefore host a huge variety of photon-mediated spin models. However, large-scale CPW lattices with non-trivial band structures have never before been successfully combined with superconducting qubits. Here we present the first such device featuring a quasi-1D CPW lattice with multiple transmon qubits. We demonstrate that superconducting-qubit readout and diagnostic techniques can be generalized to this highly multimode environment and observe the effective qubit-qubit interaction mediated by the bands of the resonator lattice. This device completes the toolkit needed to realize CPW lattices with qubits in one or two Euclidean dimensions, or negatively-curved hyperbolic space, and paves the way to driven-dissipative spin models with a large variety of connectivities.
@@ I Introduction
Quantum spin models underlie a wide variety of solid-state phenomena from ferromagnetism, to spin glasses, to parts of the phase diagram of high-temperature superconductors [girvin, SpinGlass, AndersonSpinLiquid]. Efficient classical simulation algorithms exist in special types of connectivity, but the general phenomenology of these systems remains an outstanding grand challenge, particularly when a spin Hamiltonian is accompanied by drive and dissipation. Implementing these models in synthetic quantum systems, which offer access to novel local or dynamical measurements, holds the potential to provide new insights into this challenge. Here we present an experimental platform suitable for implementing spin-spin interactions with a large variety of connectivities, spanning from simple 1D chains, to 2D, to positively-curved spherical spin models [Houck_Nature_QS, bienias:Hyperbolic], and even to hyperbolic spin models, by leveraging the flexible connectivity of microwave resonator arrays [Kollar:2019hyperbolic, Kollar:2019linegraph]. We demonstrate integration of transmon qubits without compromising the versatility of the resonator array, as well as qubit-qubit interactions mediated by the bands of the array, completing the experimental toolkit needed to realize two-dimensional or hyperbolic spin models.
@@
The crucial ingredient underlying the flexibility of this platform is the use of microwave photons to mediate interactions between effective spin-$1/2$ degrees of freedom formed by qubits. Direct qubit-qubit coupling inherently gives rise to a spin-spin interaction which follows the geometry of the device. By contrast, using a photon to mediate coupling allows the geometry of the spin-spin interaction to be decoupled from the physical locations of the qubits, and in these cases, the limits on achievable interactions are set by the ability to control and modify the propagation of the mediating photon. In the simplest case, a single-mode cavity mediates an all-to-all interaction between qubits, leading to trivial connectivity. An interaction that generates a spin model with non-trivial connectivity requires the presence of multiple modes in a bandwidth set by the qubit-photon interaction strength [Douglas_2015, bienias:Hyperbolic, Sundaresan, Benedetto2024]. In such cases, the dispersion relation and spatial profiles of the photon modes determine the interaction profile.
@@
The circuit QED platform [Blais:revmodphys], featuring superconducting qubits coupled to microwave resonators, is an ideal host for such models due to the combination of two types of hardware properties. (i) Superconducting microwave resonators have very high quality factors and can be implemented in a wide variety of different ways, with highly variable form factors on chip. As a result of this flexibility, superconducting microwave resonator arrays have been shown to host linear, quadratic, and flat bands in both one and two dimensions [Underwood:imaging, Houck:earlylattice, Houck_Nature_QS, Kollar:2019linegraph, Painter_Markov, Painter_SSH, Painter_2023, Martinez_2023, Simone:resonatorarray], and non-Euclidean structures [Kollar:2019hyperbolic]. Flat bands are particularly important since they quench kinetic energy, permitting interaction-driven strongly correlated phases to emerge in many-body systems. Furthermore, beyond their intrinsic interest, many-body systems on manifolds with spatial curvature enable experimental access to exotic topological observables. (ii) Superconducting qubits are frequency tunable, and natively exhibit strong light-matter coupling, which can easily be adjusted by orders of magnitude. The combination of these two properties gives unprecedented ability to tailor the spatial structure of strong qubit-qubit interactions, without the loss in the photon modes significantly damaging native qubit coherence times.
@@
Multimode cavity-QED models with the potential to host boson-mediated spin models with non-trivial connectivity have also been studied in atomic systems with multiple discrete modes [kollar2017supermode, Papagreorge2016, Kollar:apparatus, Vaidya:2018fp, Gopalakrishnan09, Gopalakrishnan10, Gopalakrishnan:2011jx, Gopalakrishnan:2012cf, Guo:2019ej, Guo:2019fr, Marsh2024, Vuletic:2001ur, vuletic:prl, schine:landau, schine:FloquetPolariton, schine:Laughlin]; or a continuum of modes from a waveguide [RevModPhys_Waveguides, Fayard_2021], a photonic crystal [Douglas_2015], or a BEC [Krinner:2018, Stewart:2020, Lanuza:2022]; or with Raman coupling in a single-mode cavity [Periwal_2021]. However, none of these systems can achieve the large coupling strengths and the variety of connectivities possible in superconducting microwave circuits.
@@
Previous multimode QED and waveguide QED experiments with superconducting circuits have used modulated waveguides [Liu, Sundaresan], simple one-dimensional chains of resonators [vrajitoarea2024, Simone:resonatorarray, Painter_SSH, Painter_Markov, Painter_2023], or 2D Hofstadter lattices [Owens_2018, Schuster:Chiralcavity]. In these cases, the attainable dispersion relations and spatial profiles of the photon modes have been limited by either the simplicity of the waveguide chain or by the rigid form factor of the 3D cavities. In contrast, coplanar-waveguide (CPW) resonators [Goppl_2008] can form resonator arrays [Houck_Nature_QS, Houck:earlylattice, Koch:AnnPhysBerl, Underwood:imaging] with low-disorder [Houck:earlylattice] and highly flexible connectivities [Kollar:2019hyperbolic, Kollar:2019linegraph]. It has been demonstrated that this technique can produce lattices in which the photons propagate in an effective negatively-curved hyperbolic space [Kollar:2019hyperbolic], and it has been shown theoretically that the photon-mediated qubit-qubit interaction that a hyperbolic CPW lattice would produce also follows the hyperbolic metric [bienias:Hyperbolic]. In general, CPW resonator lattices also have a strong tendency to produce flat bands that consist entirely of wave functions which flip sign on neighboring sites [Kollar:2019hyperbolic, Kollar:2019linegraph], and which can therefore give rise to short-range frustrated magnetic interaction between qubits. However, previous experimental studies of coplanar-waveguide arrays have either been limited to simple 1D chains with many qubits but large disorder [Fitzpatrick_2017], or higher-dimensional connectivity and low disorder but no qubits [Underwood:imaging, Houck:earlylattice, Kollar:2019hyperbolic].
@@
Here we present the first CPW lattice device featuring both a large-scale low-disorder resonator array with non-trivial band structure and multiple superconducting qubits. Our device features a Euclidean quasi-one-dimensional microwave resonator lattice with a non-trivial unit cell, which gives rise to conventional quadratic bands, linearly dispersing bands, and flat bands with localized eigenstates. We incorporate three flux-tunable transmon qubits directly into the CPW lattice and show that the low-disorder microwave resonator array survives the addition of the qubits and their associated flux-bias control.
@@
We demonstrate that conventional superconducting-qubit readout and diagnostic techniques [Blais:revmodphys], though originally developed for single-mode systems, can be applied in this highly multimode environment, and can be used to observe the effective qubit-qubit interaction mediated by the bands of the resonator lattice. Additionally, we extend prior non-linear spectroscopy techniques in superconducting circuits, such as Kerr spectroscopy [Bosman_2017, Peugeot_2024] and qubit two-tone spectroscopy [Blais:revmodphys], to implement reliable in-situ measurements of the mode spectrum of the CPW lattice. The technique, which we term mode-mode spectroscopy, leverages the presence of qubits to suppress the effects of background modes and does not require two-port transmission of all modes, enabling observation of localized modes as well as delocalized ones. Previous related techniques were either extremely time-consuming [Houck:earlylattice] or highly sensitive to the location of probe ports and unable to distinguish lattice modes from some types of parasitic modes induced by device packaging [Kollar:2019hyperbolic]. The device and measurement techniques presented here pave the way to CPW-lattice-based spin models with frustrated, 2D, and even hyperbolic connectivity.
@@ II Background
Traditional cavity QED describes an interaction between a single qubit/atom/spin and a single photon mode. Extensions to a continuum of photon modes have been studied in the context of emitters coupled to a waveguide, known as waveguide QED [RevModPhys_Waveguides, Fayard_2021]. CPW lattices produce a large number of discrete modes which approximate the continuous density of states of the bands of an infinite lattice. These systems therefore approximate the full continuum of waveguide QED. To simplify the language and highlight the commonalities between different ways of creating multimode systems, we refer to all systems in which an effective continuum of highly-coupled modes with a dispersion relation is relevant to forming photon-mediated interactions as (generalized) photonic crystal and waveguide QED systems. In the remainder of this section, we present a general framework for describing multimode QED devices of this kind with multiple qubits. Characteristics of these models which are specific to superconducting circuits and CPW resonators, in particular, will be discussed in Section III.
@@ II.1 Cavity and Waveguide QED
The simplest Hamiltonian for a cavity-qubit coupled system is known as the Jaynes-Cummings Hamiltonian [Kimble:2005wd, Carmichael:LesHouches, wallsMilburn]:
@@
In this work we consider systems where the waveguide is formed by a resonator network. We therefore present the general discrete-mode case of waveguide QED here, but extension to a full continuum is straightforward [cohentanoudji]. In this framework the qubit and cavity portions of the Hamiltonian become
@@
$H_{int}=\hbar g_{0}\sum_{k,j}\left(\psi^{*}_{k}(x_{j})a_{k}^{{\dagger}}\sigma_{j,-}+\mbox{ h.c.}\right)\!.$ (5)
@@ II.2 Photon Bound States and Photon-Mediated Interactions
Photonic band gaps [John_Wang] are typically created through some form of periodic modulation of the medium. One type of device architecture involves periodic modulation of the index of refraction, for the case of optical photons [Emanuel:revmodphys], or the impedance [Liu, Sundaresan], for microwave photons, and generally has a full continuum of delocalized modes. Another class of devices instead uses periodic arrays of localized sites such as optical or microwave resonators [Houck:earlylattice, Houck_Nature_QS, Kollar:2019hyperbolic, Simone:resonatorarray, Owens_2018, Schuster:Chiralcavity, Painter_Markov, Painter_SSH, Martinez_2023, Arrangoiz_Arriola_2018]. In these devices, the photonic modes are delocalized only by the coupling between the sites, and can be thought of as akin to tight-binding models of solids [girvin]. The device presented in this work is chosen to be of the latter kind because the tight localization of the mode in the individual resonators leads to separation between the hardware properties that set the on-site energy and the properties that control the coupling, or hopping, between sites.
@@
In the context of photonic band gaps, photon-mediated qubit-qubit interactions are often described as arising from the formation of photonic bound states [John_Wang, Douglas_2015, Calaj_2016]. These are localized modes with photonic character which emerge even if all of the photon modes are delocalized in the absence of coupling to the qubit. If the qubit is in the band gap, there is an intuitive picture of the formation of the bound state in which the qubit hybridizes with the photon modes and acquires a cloud of virtual photons surrounding it. In the conventional case, where the dispersion of the photon modes above the band edge is quadratic ($\omega_{k}=\omega_{edge}+\beta k^{2}$) and the coupling constants $g_{j,k}$ are approximately equal, this photon cloud is known to take the form of an exponentially localized envelope, $e^{-|x|/L}$ [John_Wang, Douglas_2015, Calaj_2016], where $|x|$ is the distance from the qubit, and where the localization length $L$ is determined by both the band curvature $\beta$, and the detuning between the qubit and the band edge. The simple picture of the photonic bound state being a cloud of virtual photons surrounding a qubit is valid only for a qubit well inside the band gap. However, an exact solution of similar form exists at any detuning from an ideal quadratic band edge and has been derived in Refs. [John_Wang, Liu, Sundaresan, Calaj_2016]. If two qubits are coupled to the band edge, then interactions between them can be thought of as arising from the tail of the photonic bound state localized around one qubit driving transitions in the other qubit. Hence, the spatial structure of the photonic bound state dictates the spatial structure of the interaction.
@@
Alternatively, the photon-mediated qubit-qubit interactions can be modeled directly as a photon-exchange process [Douglas_2015, Sundaresan, bienias:Hyperbolic], as described in Section II.1. For arbitrary detunings, the frequency of the photonic bound state and the resulting interaction between the $j^{th}$ qubit and the $l^{th}$ qubit must be found self-consistently [Sundaresan, bienias:Hyperbolic, Calaj_2016]. However, if both qubits are in the band gap, then second-order perturbation theory can be used to obtain a simple form valid for any dispersion
@@
where $\sigma_{j,-}$ and $\sigma_{j,+}$ indicate lowering and raising operators on the $j^{th}$ qubit, and $\Delta_{k}$ is the detuning between the qubits and mode $k$ (assumed to be approximately equal for the two qubits). A quadratic band edge in a one-dimensional set of modes gives a simple exponential interaction, but this can be changed by going to higher dimensions, non-quadratic dispersions [Yidan:2022], or non-plane-wave modes.
@@ III Photonic Crystals in Circuit QED
Since the form of the qubit-qubit interaction mediated by a photonic bound state is determined by the dispersion and mode functions of the band, creating spin models with a large variety of connectivities requires the ability to engineer diverse band structures with strong qubit-photon coupling. In this section we describe the hardware aspects of CPW lattices that make them ideally suited to this task, as well as the measurement techniques native to this microwave-frequency implementation of waveguide QED. We emphasize that the methods developed for single-mode circuit QED [Blais:revmodphys] can be generalized to a full toolkit for design, characterization, and measurement of multimode waveguide-QED devices made with CPW resonators.